OpenStudy (anonymous):

What is the distance between points (2, 8) and (7, 5)? Round to the nearest tenth of a unit.

OpenStudy (anonymous):

how do i do this

OpenStudy (anonymous):

\[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] distance formula your points are the ones given.

OpenStudy (anonymous):

i got 4?

OpenStudy (anonymous):

@iambatman

TheSmartOne (thesmartone):

4 is wrong.

TheSmartOne (thesmartone):

\(\sf (x_1,y_1),(x_2,y_2)\) \(\sf (2,8)~~~,(7,5)\) Plug them in to the formula that batman gave you :)

OpenStudy (anonymous):

he gave me the wrong formula?

OpenStudy (anonymous):

he said x2-x1 not x1-y1

OpenStudy (anonymous):

oh nvm

TheSmartOne (thesmartone):

The points you have been given is \(\sf (2,8)\) and that is in the form of \(\sf (x_1,y_1)\) So \(\sf x_1 =2 \\ y_1 =8\)

OpenStudy (anonymous):

i got 22

OpenStudy (anonymous):

then 4.7

OpenStudy (anonymous):

is this right

TheSmartOne (thesmartone):

What is your final answer?

TheSmartOne (thesmartone):

If we plug in the numbers we get... \(\sf\begin{aligned}\sqrt{\left( 7 - 2 \right)^2 + \left( 5 - 8 \right)^2} \end{aligned}\)

TheSmartOne (thesmartone):

Solve from there.

OpenStudy (freckles):

\[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}= \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] @iambatman gave you correct formula

OpenStudy (anonymous):

i got 4.7

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

5.8?

OpenStudy (anonymous):

@Michele_Laino @MCLover1477 i need help

OpenStudy (michele_laino):

we have to compute this: \[d = \sqrt {{{\left( {2 - 7} \right)}^2} + {{\left( {8 - 5} \right)}^2}} = ...?\]

OpenStudy (anonymous):

this is kinda easy um well so do you have multiple choice answers or do you have to put a number?

OpenStudy (anonymous):

@alcathcart ?

OpenStudy (anonymous):

we type an answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

25+9

OpenStudy (anonymous):

oh i think i forgot that negative times a negative is positive

OpenStudy (michele_laino):

ok! what is sqrt(25+9)=...?

OpenStudy (michele_laino):

\[d = \sqrt {{{\left( {2 - 7} \right)}^2} + {{\left( {8 - 5} \right)}^2}} = \sqrt {25 + 9} = ...?\]

OpenStudy (anonymous):

5.8

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

wow thank you!! can you help with the rest just a few more